TANGENTS AND NORMALS
find the points on the curve  2a^2y=x^3 – 3ax^2 where the tangent is parallel to x axis 

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

2a^2y = x^3 - 3ax^2
Slope = 0
2a^2y' = 3x^2 - 6ax
3x^2 - 6ax = 0
3x(x-2a) = 0
x = 0, 2a
Put in curve:
x = 0\Rightarrow y = 0
x = 2a\Rightarrow y = -2a